# Synthetic bootstrapped pretraining

### v2

@@ 4 Experiment results
We perform SBP experiments under the compute-matched framework outlined in §3 at two levels of training compute budget: 200B-scale and 1T-scale. After joint training on real and synthetic data $\{{{\mathcal{D}}_{\text{pretrain}}},{{\mathcal{S}}_{\text{pretrain}}}\}$, we find SBP consistently improves upon the repetition baseline throughout both scales (Table 1). In this section, we focus on presenting the performance of SBP and evaluating the quality of the synthesized pretraining data. We defer the implementation details of SBP to §A.1.
Table 1: Computed-matched comparison of Synthetic Bootstrapped Pretraining (SBP) and oracle performance gains over the repetition baseline. On average, SBP delivers roughly 43% of the performance improvement in QA accuracy attainable by an oracle with access to 20x more unique data.
200B-scale 1T-scale
Benchmark Baseline SBP Oracle Baseline SBP Oracle
@@
OpenWebText2 5.74 -0.53 -1.02 4.51 -0.02 -0.12
LAMBADA 6.87 -0.85 -1.86 4.33 -0.03 -0.22
Five-shot MMLU 3.83 -0.36 -0.51 3.17 -0.06 -0.05
@@
ARC-Challenge (0-shot) 35.32 +1.28 +2.82 42.66 +1.62 +3.84
ARC-Easy (0-shot) 68.94 +2.65 +4.29 75.63 +0.42 +2.11
SciQ (0-shot) 90.50 +1.00 +2.40 93.20 +0.80 +0.50
Winogrande (0-shot) 60.14 +1.90 +5.53 65.19 +1.42 +2.92
TriviaQA (1-shot) 22.51 +3.36 +7.37 36.07 +0.25 +0.59
WebQS (1-shot) 8.56 +3.74 +10.83 19.34 +0.54 +0.44
Average QA accuracy 47.66 +2.32 +5.54 55.35 +0.84 +1.73
@@ 4.1 Main benchmark performance
The training dynamics of SBP partly reveal its core mechanism. As we can see in Figure 2, initially, the baseline performs similarly to the oracle, since their training data share the same distribution, and when the number of tokens seen is small, there is no distinction between the two. Then gradually, the oracle becomes a better model than the baseline, as it has access to unlimited unique training data. For the SBP dynamics, it initially performs worse than both the baseline and the oracle, which is expected since the quality of the synthesized data at most matches that of the real data. However, gradually, the SBP continues to scale while the baseline has plateaued. This suggests that ${\mathcal{S}}_{\text{pretrain}}$} offers a signal ${\mathcal{D}}_{\text{pretrain}}$} alone cannot capture.
Lastly, to validate the benefit of SBP across different training scales, we implement a larger experiment with $\|{{\mathcal{D}}_{\text{pretrain}}}\|=$50B unique tokens under a compute budget of 1T total training tokens. We generate 125B total synthetic tokens $\|{{\mathcal{S}}_{\text{pretrain}}}\|=$125B, with similar ablation presented in §A.2, and adhere to the same no-repetition-for-synthetic-data principle adopted at the 200B-scale. Examining the oracle improvement from Table 1, we can see that the perplexity-based measurements and most QA benchmarks have plateaued at this scale (Liu et al., 2023). However, ARC-Challenge and Winogrande continue to deliver smooth performance changes, making them suitable candidates for tracking pretraining capability at large scales. In particular, in ARC-Challenge, both SBP and Oracle yield a larger performance improvement than their 200B-scale counterparts. That said, other benchmarks still provide a directional signal of capability improvement. This demonstrates the advantage of having a diverse collection of evaluation benchmarks covering a wide range of difficulties. On average, SBP delivers 0.84/1.73=0.84/1.73=48% of the improvement in QA accuracy attained by the oracle.
@@ 4.2 Analysis of synthetic data
We start by showing some samples of synthesized documents from the 200B-scale experiment (Figure 3) with more samples from 1T-scale presented in §B.4. On the left, we display a real document about a practical, first-person guide to the coffee houses in San Diego. Then, we present two synthesized texts that exhibit notable differences in both framing and depth, with varying degrees of fidelity to the seed document. Synthesis I sticks to the same topic but shifts toward an expository essay on espresso machines and bean quality, with little mention of specific coffee shops. Synthesis II adopts a promotional, comparative style, linking San Diego’s coffee culture to New York’s and praising Café Lestat in a way that departs from the original’s balanced assessments. Note that SBP provides no instructions on how the synthesizer should use the seed texts to write new documents. The model spontaneously learns to introduce new content and style into the discussion while staying on topic. In fact, it is challenging to manually craft a prompt to an instruction-tuned model that would output either Synthesis I or II with the real document as input. This example highlights how SBP differs from existing paradigms of data synthesis – the output somehow first abstracts the seed document and then synthesizes new text with more generalized narratives, genres, and intent. We provide more extensive analysis of this observation in §5.
@@
Non-factual failure stems from hallucinations that introduce non-existent entities or relations inconsistent with reality. We find that synthesis at the 1T-scale significantly reduces these errors compared to the 200B-scale. As the data synthesizer is trained on more data, the factuality of the generated outputs tends to converge toward that of real data.
@@ 5.1 A hierarchical concept model for natural language
The additional signal from the posterior then enables a form of self-distillation. The synthesizer, by learning a more complex conditional objective, becomes a more knowledgeable “teacher” model that has learned to infer the latent structure of data. The synthetic data it produces is then the knowledge “distilled” from this teacher (Hinton et al., 2015). The final LM training then acts as a “student” that learns from a combination of real and synthetic data, allowing it to discover information that real data alone cannot reveal.
@@ 5.2 From idealized models to language model reality
For real text documents, we do not know the true data-generating process, and any parametric assumption would be incorrect. This is where the power of the transformer neural network shines. A transformer is a mapping-first (Breiman, 2001) approach. It does not require explicit modeling of the underlying parametric model. Instead, as a universal function approximator (Candès, 1998), it directly learns the complex conditional distribution $p_{\theta}({d_{2}}|{d_{1}})$ from paired data alone.
@@ 6 Discussion
Before the prevalence of pretraining (Radford et al., 2018b; Devlin et al., 2019), we would need 40M pairs of English and French sentences (Sutskever et al., 2014) to grant LM the ability to translate from one language to another. In contrast, any modern LM (Gemini, 2024; OpenAI et al., 2024; Gunter et al., 2024) can easily achieve this task via a single prompt. Such impressive advancement owes its origin to the rich correlations between words within a document LM has learned during pretraining. This shift from a hard-to-collect, task-specific dataset to the readily available, generically crawled internet data marks a transition from relying on strong but scarce supervision to a weak self-supervision that is easy to scale. As we gradually deplete this weak source of self-supervision by exhausting the available internet data, many have called for stronger forms of supervision, such as reinforcement learning (DeepSeek-AI et al., 2025; OpenAI, 2024). We instead offer an alternative perspective that continues to search for a form of self-supervision weaker than next-token prediction. SBP offers a particular instantiation of such an effort by examining the correlations between documents missed by the current pretraining paradigm. It remains to explore other forms of supervision not currently utilized.
The fact that SBP could provide any improvement originates from the poor inductive bias of the transformer neural network. For example, transformers trained on the text “A is B” can not generalize to “B is A” (Berglund et al., 2023), requiring the user to curate synthetic data that explicitly narrates the converse relation (Yang et al., 2025). Similarly, one can imagine an architecture with very good inductive bias such that the LM trained individually on $d_{1}$} and $d_{2}$} can automatically internalize the relation between the two. Despite such poor inductive bias, transformers have the impressive advantage of being more parallelizable and scalable than their alternatives (Vaswani et al., 2017; Shazeer et al., 2017). Given this trade-off, SBP offers a unique solution that preserves the system benefits of the transformer architecture while also enabling the learning of missed correlations by encoding such additional signals via synthetic data.
@@ 6.1 Future directions
In our implementation of SBP, we use Qwen3-0.6B-Embedding (Zhang et al., 2025) to obtain embeddings of DCLM (Li et al., 2024) documents. An ideal implementation of SBP would only rely on the 3B-parameter model and the pretraining dataset itself to curate the paired synthesizer-tuning dataset. To achieve this, we can use the activations of the self-attention layer from an intermediate transformer block as a learned representation of documents. Khandelwal et al. (2020) and Yang et al. (2023) implemented this at the much smaller scale of $\sim 300$M parameters and $\sim 3$B tokens. However, our experiments operate at a much larger scale with a customized model. As a result, we utilize the optimized vLLM (Kwon et al., 2023) inference infrastructure for Qwen3-0.6B embedding models to efficiently index the pretraining corpus. Since the SBP procedure only requires a coarse binary decision of relevant vs. not relevant, which is much weaker than fine-grained document ranking embedding models are optimized for, we leave the more involved inference infrastructure for future work.
@@
One experimental consequence of LM pretraining is a clean scaling law (Kaplan et al., 2020, Equation 1.4) that relates the held-out test loss $L(N,D)$ to the number of LM parameters $N$ and the size of the pretraining dataset $D$. A natural scientific question is how the scaling law of SBP compares to the scaling law of pretraining. In our experiments, we essentially evaluate $L(N,D)$ with N=3N=3B at two different points D=D=10B and DD=50B. There are two obstacles to a full scaling law: first, SBP is inherently a large-scale algorithm that cannot be scaled down. Since SBP first uses the pretrained LM to generate synthetic text and then trains on it, if the model and dataset sizes are too small, the generated text may not even be coherent. In contrast, the scaling experiments in Kaplan et al. (2020) involve model sizes ranging from 768M to 1.5B and dataset sizes ranging from 22M to 23B, which allows for efficient experimentation. Second, varying $N$ or $D$ implies redoing the synthesizer-tuning and subsequent data synthesis over billions of tokens. Additionally, varying $D$ also implies redoing the nearest neighbor matching, as the neighbors are only defined given a fixed document pool. Obstacles aside, it would be interesting to see whether the SBP scaling law differs from the normal scaling law by a smaller multiplicative factor or a better exponent. Since SBP additionally utilizes inter-document correlations, a form of long-range interactions, its scaling law not only helps us understand SBP but also potentially helps us better understand natural language data itself (Ebeling & Pöschel, 1994).
@@ 6.2 Conclusion
In conclusion, we introduce Synthetic Bootstrapped Pretraining (SBP) as a new LM pretraining framework that leverages inter-document correlations missed by the standard pretraining objective. We demonstrate the effectiveness of SBP by pretraining a 3B-parameter model from scratch for up to 1T tokens under a rigorously designed compute-matched experiment setup. Qualitative and quantitative analyses of the synthesized text reveal rich variation that cannot be captured by simple paraphrases. In addition to being practically effective, SBP admits a natural Bayesian interpretation, where it implicitly learns a posterior that infers the latent concept in a given document. Its bootstrapping nature grants SBP the possibility of scaling the pretraining capability of the LM beyond its current limit.
@@ A.1 SBP implementation details
Recall from §3 that we work with a 3B-parameter transformer architecture and pretraining dataset at $\|{{\mathcal{D}}_{\text{pretrain}}}\|=$10B and $\|{{\mathcal{D}}_{\text{pretrain}}}\|=$50B scale. To take advantage of efficient ANN search at pretraining scale, we embed the documents from ${\mathcal{D}}_{\text{pretrain}}$} as 1,024 dimensional vectors using Qwen3-Embedding-0.6B. Then, we use ScaNN (Guo et al., 2020; Sun et al., 2023) with 8-bit quantization to perform efficient similarity search. We adopt an asymmetric sharding to keys and value vectors. For each value vector, we build a ScaNN search tree with $\sqrt{N}$ leaves where $N$ is the number of vectors in each value shard. To distribute the key shards across each search tree, we employ a “salting” strategy, where we create multiple copies of the ScaNN searcher and assign one key shard to each salted copy of the searcher (Figure 4). This design enables us to perform a top-200 nearest neighbor search over $|{{\mathcal{D}}_{\text{pretrain}}}|=$60M documents within 155M CPU hours.
@@
At both the 200B-scale and 1T-scale, after obtaining the top 200 neighbors for each sample, we select the pairs whose similarity score is greater than 0.75. We chose this cut-off as it would later lead to a tractable size of synthesizer-tuning dataset ${\mathcal{D}}_{\text{ST}}$}. To access the effect of choosing a different threshold, we provide a quantitative analysis of the fraction of relevant documents around each bin of similarity threshold in Figure 5 using the same metric defined in §4.2. We can see that a larger similarity score yields pairs with higher relevance but also more duplicates. Finally, we eliminate near-duplicates using a rule-based filtering approach. The dedup process involves first normalizing text by removing punctuation, converting to lowercase, and eliminating numbers, followed by tokenization using SentencePiece. We then generate “shingles” using 13-token sliding windows within $d_{1}$}. Training pairs are discarded if any shingle from $d_{1}$} appears in $d_{2}$}.
@@
Finally, we perform the hierarchical sampling procedure defined in §2 with a temperature of 1.0 and top_p threshold 0.9. We apply a rule-based filtering that removes synthesized documents containing repeated occurrences of 13-token shingles. This effectively removes texts with repetition failure. We use vLLM (Kwon et al., 2023) and obtain a throughput of 8.3K tokens per B200 second. This amounts to 2.5K B200 hours for the 200B-scale synthesis and 4.2K B200 hours for the 1T-scale synthesis.

### v3

@@ 4 Experiment results
We perform SBP experiments under the compute-matched framework outlined in §3 at three levels of training compute budget: 200B-scale, 1T-scale (3B), and 1T-scale (6B). After joint training on real and synthetic data $\{{{\mathcal{D}}_{\text{pretrain}}},{{\mathcal{S}}_{\text{pretrain}}}\}$, we find SBP consistently improves upon the repetition baseline across all scales (Table 1). In this section, we focus on presenting the performance of SBP and evaluating the quality of the synthesized pretraining data. We defer the implementation details of SBP to §A.1.
Table 1: Computed-matched comparison of Synthetic Bootstrapped Pretraining (SBP) and oracle performance gains over the repetition baseline. On average, SBP delivers roughly 43% of the performance improvement in QA accuracy for the 3B model and 58% for the 6B model, attainable by an oracle with access to 20x more unique data.
200B-scale 1T-scale (3B) 1T-scale (6B)
Benchmark Baseline SBP Oracle Baseline SBP Oracle Baseline SBP Oracle
@@
OpenWebText2 5.74 -0.53 -1.02 4.51 -0.02 -0.12 4.25 -0.06 -0.21
LAMBADA 6.87 -0.85 -1.86 4.33 -0.03 -0.22 3.63 -0.06 -0.25
Five-shot MMLU 3.83 -0.36 -0.51 3.17 -0.06 -0.05 3.08 -0.08 -0.13
@@
ARC-Challenge (0-shot) 35.32 +1.28 +2.82 42.66 +1.62 +3.84 47.44 +0.77 +0.17
ARC-Easy (0-shot) 68.94 +2.65 +4.29 75.63 +0.42 +2.11 78.70 +0.51 +0.85
SciQ (0-shot) 90.50 +1.00 +2.40 93.20 +0.80 +0.50 92.90 +1.90 +1.80
Winogrande (0-shot) 60.14 +1.90 +5.53 65.19 +1.42 +2.92 70.17 +0.47 +2.36
TriviaQA (1-shot) 22.51 +3.36 +7.37 36.07 +0.25 +0.59 40.64 +0.49 +3.19
WebQS (1-shot) 8.56 +3.74 +10.83 19.34 +0.54 +0.44 19.88 +3.79 +5.22
Average QA accuracy 47.66 +2.32 +5.54 55.35 +0.84 +1.73 58.29 +1.32 +2.26
@@ 4.1 Main benchmark performance
The training dynamics of SBP partly reveal its core mechanism. As we can see in Figure 2, initially, the baseline performs similarly to the oracle, since their training data share the same distribution, and when the number of tokens seen is small, there is no distinction between the two. Then gradually, the oracle becomes a better model than the baseline, as it has access to unlimited unique training data. For the SBP dynamics, it initially performs worse than both the baseline and the oracle, which is expected since the quality of the synthesized data at most matches that of the real data. However, gradually, SBP continues to scale while the baseline has plateaued. This suggests that ${\mathcal{S}}_{\text{pretrain}}$} offers a signal ${\mathcal{D}}_{\text{pretrain}}$} alone cannot capture.
Lastly, to validate the benefit of SBP across different training scales, we implement a larger experiment with $\|{{\mathcal{D}}_{\text{pretrain}}}\|=$50B unique tokens under a compute budget of 1T total training tokens using both 3B and 6B-parameter models. Based on the ablation studies presented in §A.2, we include $\|{{\mathcal{S}}_{\text{pretrain}}}\|=$125B synthetic tokens for the 3B model and $\|{{\mathcal{S}}_{\text{pretrain}}}\|=$250B synthetic tokens for the 6B model, adhering to the principle of no repetition for synthetic data. Examining the results in Table 1, we observe that while perplexity-based measurements plateau for the 3B model (samepretrain), benchmarks like ARC-Challenge and Winogrande continue to show significant gains. On average, SBP recovers 0.84/1.73=0.84/1.73=48% of the oracle’s QA accuracy improvement for the 3B model. The results are even more pronounced for the 6B model, where SBP delivers a relative improvement of $1.32/2.26=58\%$ compared to the oracle. This demonstrates that SBP’s effectiveness scales favorably with model size. Furthermore, the increased optimal synthetic data ratio for the 6B model suggests that larger models possess greater capacity to exploit the additional information encoded in the synthetic corpus.
@@ 4.2 Analysis of synthetic data
We first show some samples of synthesized documents from the 200B-scale experiment (Figure 3) with more samples from 1T-scale (3B) presented in §B.4. On the left, we display a real document about a practical, first-person guide to the coffee houses in San Diego. Then, we present two synthesized texts that exhibit notable differences in both framing and depth, with varying degrees of fidelity to the seed document. Synthesis I sticks to the same topic but shifts toward an expository essay on espresso machines and bean quality, with little mention of specific coffee shops. Synthesis II adopts a promotional, comparative style, linking San Diego’s coffee culture to New York’s and praising Café Lestat in a way that departs from the original’s balanced assessments. Note that SBP provides no instructions on how the synthesizer should use the seed texts to write new documents. The model spontaneously learns to introduce new content and style into the discussion while staying on topic. In fact, it is challenging to manually craft a prompt to an instruction-tuned model that would output either Synthesis I or II with the real document as input. This example highlights how SBP differs from existing paradigms of data synthesis – the output somehow first abstracts the seed document and then synthesizes new text with more generalized narratives, genres, and intent. We provide more extensive analysis of this observation in §5.
@@
Non-factual failure stems from hallucinations that introduce non-existent entities or relations inconsistent with reality. We find that synthesis at the 1T-scale (3B) significantly reduces these errors compared to the 200B-scale. Furthermore, with the 6B-parameter model, the non-factual rate further decreases to 6.5%. This indicates that as the data synthesizer is trained on more data and with larger models, the factuality of the generated outputs tends to converge toward that of real data.
@@ 5.1 A hierarchical concept model for natural language
The additional signal from the posterior then enables a form of self-distillation. The synthesizer, by learning a more complex conditional objective, becomes a more knowledgeable “teacher” model that has learned to infer the latent structure of data. The synthetic data it produces is then the knowledge “distilled” from this teacher (hinton2015distillingknowledgeneuralnetwork). The final LM training then acts as a “student” that learns from a combination of real and synthetic data, allowing it to discover information that real data alone cannot reveal.
@@ 5.2 From idealized models to language model reality
For real text documents, we do not know the true data-generating process, and any parametric assumption would be incorrect. This is where the power of the transformer neural network shines. A transformer is a mapping-first (2cultures) approach. It does not require explicit modeling of the underlying parametric model. Instead, as a universal function approximator (emmanuelthesis), it directly learns the complex conditional distribution $p_{\theta}({d_{2}}|{d_{1}})$ from paired data alone.
@@ 6 Discussion
Before the prevalence of pretraining (gpt2; devlin-etal-2019-bert), we would need 40M pairs of English and French sentences (sutskever2014sequencesequencelearningneural) to grant LM the ability to translate from one language to another. In contrast, any modern LM (gemini; gpt4; gunter2024appleintelligencefoundationlanguage) can easily achieve this task via a single prompt. Such impressive advancement owes its origin to the rich correlations between words within a document LMs have learned during pretraining. This shift from a hard-to-collect, task-specific dataset to the readily available, generically crawled internet data marks a transition from relying on strong but scarce supervision to a weak self-supervision that is easy to scale. As we gradually deplete this weak source of self-supervision by exhausting the available internet data, many have called for stronger forms of supervision, such as reinforcement learning (r1; o1). We instead offer an alternative perspective that continues to search for a form of self-supervision weaker than next-token prediction. SBP offers a particular instantiation of such an effort by examining the correlations between documents missed by the current pretraining paradigm. It remains to explore other forms of supervision not currently utilized.
The fact that SBP could provide any improvement originates from the poor inductive bias of the transformer neural network. For example, transformers trained on the text “A is B” can not generalize to “B is A” (berglund2023reversal), requiring the user to curate synthetic data that explicitly narrates the converse relation (entigraph). Similarly, one can imagine an architecture with very good inductive bias such that the LM trained individually on $d_{1}$} and $d_{2}$} can automatically internalize the relation between the two. Despite such poor inductive bias, transformers have the impressive advantage of being more parallelizable and scalable than their alternatives (transformer; shazeer2017outrageouslylargeneuralnetworks). Given this trade-off, SBP offers a unique solution that preserves the system benefits of the transformer architecture while also enabling the learning of missed correlations by encoding such additional signals via synthetic data.
@@ 6.1 Future directions
In our implementation of SBP, we use Qwen3-0.6B-Embedding (zhang2025qwen3embeddingadvancingtext) to obtain embeddings of DCLM (li2024datacomplm) documents. An ideal implementation of SBP would only rely on the 3B-parameter model and the pretraining dataset itself to curate the paired synthesizer-tuning dataset. To achieve this, we can use the activations of the self-attention layer from an intermediate transformer block as a learned representation of documents. Khandelwal:2020 and resmem implemented this at the much smaller scale of $\sim 300$M parameters and $\sim 3$B tokens. However, our experiments operate at a much larger scale with a customized model. As a result, we utilize the optimized vLLM (kwon2023efficient) inference infrastructure for Qwen3-0.6B embedding models to efficiently index the pretraining corpus. Since the SBP procedure only requires a coarse binary decision of relevant vs. not relevant, which is much weaker than fine-grained document ranking embedding models are optimized for, we leave the more involved inference infrastructure for future work.
@@
One experimental consequence of LM pretraining is a clean scaling law (kaplan2020scalinglawsneurallanguage, Equation 1.4) that relates the held-out test loss $L(N,D)$ to the number of LM parameters $N$ and the size of the pretraining dataset $D$. A natural scientific question is how the scaling law of SBP compares to the scaling law of pretraining. In our experiments, we evaluate $L(N,D)$ at three distinct points: $(N=3\text{B},D=10\text{B})$, $(N=3\text{B},D=50\text{B})$, and $(N=6\text{B},D=50\text{B})$. We observe that SBP delivers larger relative improvements with the 6B model, suggesting a favorable scaling behavior. There are two obstacles to a full scaling law: first, SBP is inherently a large-scale algorithm that cannot be scaled down. Since SBP first uses the pretrained LM to generate synthetic text and then trains on it, if the model and dataset sizes are too small, the generated text may not even be coherent. In contrast, the scaling experiments in kaplan2020scalinglawsneurallanguage involve model sizes ranging from 768M to 1.5B and dataset sizes ranging from 22M to 23B, which allows for efficient experimentation. Second, varying $N$ or $D$ implies redoing the synthesizer-tuning and subsequent data synthesis over billions of tokens. Additionally, varying $D$ also implies redoing the nearest neighbor matching, as the neighbors are only defined given a fixed document pool. Obstacles aside, it would be interesting to see whether the SBP scaling law differs from the normal scaling law by a smaller multiplicative factor or a better exponent. Since SBP additionally utilizes inter-document correlations, a form of long-range interactions, its scaling law not only helps us understand SBP but also potentially helps us better understand natural language data itself (Ebeling_1994).
@@ 6.2 Conclusion
In conclusion, we introduce Synthetic Bootstrapped Pretraining (SBP) as a new LM pretraining framework that leverages inter-document correlations missed by the standard pretraining objective. We demonstrate the effectiveness of SBP by pretraining 3B and 6B-parameter models from scratch for up to 1T tokens under a rigorously designed compute-matched experiment setup. Qualitative and quantitative analyses of the synthesized text reveal rich variation that cannot be captured by simple paraphrases. In addition to being practically effective, SBP admits a natural Bayesian interpretation, where it implicitly learns a posterior that infers the latent concept in a given document. Its bootstrapping nature grants SBP the possibility of scaling the pretraining capability of the LM beyond its current limit.
@@ A.1 SBP implementation details
Recall from §3 that we work with 3B and 6B-parameter transformer architectures and pretraining dataset at $\|{{\mathcal{D}}_{\text{pretrain}}}\|=$10B and $\|{{\mathcal{D}}_{\text{pretrain}}}\|=$50B scale. To take advantage of efficient ANN search at pretraining scale, we embed the documents from ${\mathcal{D}}_{\text{pretrain}}$} as 1,024 dimensional vectors using Qwen3-Embedding-0.6B. Then, we use ScaNN (guo2020accelerating; sun2023soar) with 8-bit quantization to perform efficient similarity search. We adopt an asymmetric sharding to keys and value vectors. For each value vector, we build a ScaNN search tree with $\sqrt{N}$ leaves where $N$ is the number of vectors in each value shard. To distribute the key shards across each search tree, we employ a “salting” strategy, where we create multiple copies of the ScaNN searcher and assign one key shard to each salted copy of the searcher (Figure 4). This design enables us to perform a top-200 nearest neighbor search over $|{{\mathcal{D}}_{\text{pretrain}}}|=$60M documents within 155M CPU hours.
@@
At all scales, after obtaining the top 200 neighbors for each sample, we select the pairs whose similarity score is greater than 0.75. We chose this cut-off as it would later lead to a tractable size of synthesizer-tuning dataset ${\mathcal{D}}_{\text{ST}}$}. To access the effect of choosing a different threshold, we provide a quantitative analysis of the fraction of relevant documents around each bin of similarity threshold in Figure 5 using the same metric defined in §4.2. We can see that a larger similarity score yields pairs with higher relevance but also more duplicates. Finally, we eliminate near-duplicates using a rule-based filtering approach. The dedup process involves first normalizing text by removing punctuation, converting to lowercase, and eliminating numbers, followed by tokenization using SentencePiece. We then generate “shingles” using 13-token sliding windows within $d_{1}$}. Training pairs are discarded if any shingle from $d_{1}$} appears in $d_{2}$}.
@@
Finally, we perform the hierarchical sampling procedure defined in §2 with a temperature of 1.0 and top_p threshold 0.9. We apply a rule-based filtering that removes synthesized documents containing repeated occurrences of 13-token shingles. This effectively removes texts with repetition failure. We use vLLM (kwon2023efficient) and obtain a throughput of 8.3K tokens per B200 second. This amounts to 2.5K B200 hours for the 200B-scale synthesis, 4.2K B200 hours for the 1T-scale (3B) synthesis, and 8.4K B200 hours for the 1T-scale (6B).